find the slope,if possible, of the line passing through each pair of points (2, -1) and (6, 3)
Answer by stanbon(75887)

find the slope,if possible, of the line passing through each pair of points (2, -1) and (6, 3)
---
m = (3--1)/(6-2) = 4/4 = 1
-----
Cheers,
Stan H.